the area of a rectangle is three times more than twice its width. if the area is 35. Find the dimensions of the rectangle.

Answer by drj(1380)

The area of a rectangle is three times more than twice its width. If the area is 35. Find the dimensions of the rectangle.

Step 1. Let w be the width and l be the length.

Step 2. Area A = length * width = l*w = 35.

Step 3. A = 3*2w =35 or 6w=35 or w=35/6 and l=35/(35/6)=6

Step 4. ANSWER. The length is 6 and the width is 35/6.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J